OpenStudy (anonymous):

Please help! I'll give you a medal+fan. Estimate the diameter of a circle with a circumference of 54 feet. A. 17.2 Feet. B.34.4 Feet C.169.56 Feet D.84.78 Feet.

OpenStudy (anonymous):

A

OpenStudy (mathstudent55):

The circumference of a circle is \(C = \pi d\) That means the diameter, d, is \(d = \dfrac{C}{\pi} \) \(\pi\) is approx 3.14, so just use 3. What is 54/3 = ? That is an approximation of the diameter. Find the closest choice.

OpenStudy (anonymous):

THANK YOU!!!!